# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@ II.1.2 Self-consistent configuration recovery
$\begin{split}&P_{\textrm{recovery}}(M,N,b,\varepsilon)\geq\sum_{g=0}^{M-N-b}\sum_{h=0}^{N-b}\frac{\binom{M-N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(N-b+j)}\right)\\ &+\sum_{g=0}^{N-b}\sum_{h=0}^{M-N-b}\frac{\binom{N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(M-N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(M-N-b+j)}\right)\equiv F(M,N,b,\epsilon).\end{split}$ (S24)
Figure S2: Lower bound to the probability of recovering a particular bitstring with Hamming distance $2b$ from a reference approximation ${\bf{x}}_{r}$ bistring from a bistring sampled from the uniform distribution. The lower bound is shown as a function of the number of qubits. Different values of $b$ are shown, as indicated by the colorbar. The dashed black curve shows $1/2^{M}$, which is the probability of obtaining the desired bitsting directly from the uniform distribution. The circles connected by lines show the fraction of configurations for different families of bitstrings (as indicated by the colorbar) obtained by running the configuration recovery procedure on the samples from the quantum device on the N2 molecule for the 6-31-G and cc-pVDZ bases sets. The inset shows the lower bound as a function of $b$ for $M=52$ qubits.
II.1.3 Optimization of the circuit parameters
The optimization of circuit parameters to produce better distributions in the space of electronic configurations is an avenue of improvement of the HPC quantum estimator framework. We develop a formalism for parameter optimization using the subspace energy as the cost function, in contrast to previous experiments that rely on standard VQE optimization frameworks [30, 31]. The HPC quantum estimator is constructed by running a diagonalization procedure in $K$ subspaces defined by different batches of sampled configurations, as described in Sec. II.1.1. The number of subspaces considered $K$ can go from $K=1$ to $K=K_{\textrm{max}}$, where $K_{\textrm{max}}=\binom{D}{d}$. Recall that $D$ is the dimensionality of the subspace of the Fock space spanned by configurations with the correct particle number and $d$ is the dimension of the SCI subspace. Note that $K_{\textrm{max}}$ is double-combinatorial in the number of spin-orbitals and electrons. Therefore, computing the HPC quantum estimator (with or without configuration recovery) for all possible sets of $d$ configurations is not efficient. Instead, we propose the use of a Monte Carlo estimator over batches of configurations $\mathcal{S}^{(k)}$ to optimize the energy of the HPC quantum estimator averaged over different sets of configurations.
where $E^{(k)}$ is defined in Sec. II.1.1, Eq. (S11) as the SCI ground-state energy for the configurations in $\mathcal{S}^{(k)}$. An estimator to the double combinatorially-large summation in $\mathcal{E}(\theta)$ can be obtained by the Monte-Carlo unbiased estimator:
$\mathcal{E}(\theta)\approx\frac{1}{K}\sum_{k=1}^{K}E^{(k)}$ (S27)
whose Monte-Carlo estimator is given by:
Another practical issue with the gradient-based optimization of $\mathcal{E}(\theta)$ is the requirement to evaluate $\left\langle{\bf{x}}_{(i)}|\Psi_{\theta}\right\rangle$ and $\left\langle\partial_{\theta}\Psi_{\theta}|{\bf{x}}_{(i)}\right\rangle$, which rely in the Hadamard test for their evaluation. The implementation of the Hadamard test requires controlled versions of the circuits that realize the variational states. The implementation of such controlled unitaries requires circuit depths beyond the reach of current quantum devices.
An alternative approach is to optimize $\mathcal{E}(\theta)$ by explicit-gradient-free methods like COBYLA [69] or simulated annealing. Gradient-free optimization of the circuit parameters is used in the numerical study of Sec. III.1.4.
II.1.4 Orbital optimization
We advise the reader to first visit Sec. II.1.3, since the cost function that is optimized is similar in this scenario. Orbital or single-particle basis rotations are a common approach to improve the accuracy of variational calculations of electronic structure systems. Orbital optimizations can be applied in conjunction with a wide variety of electronic structure methods, including complete active space diagonalizations [70, 71, 72, 73], the density matrix renormalization group [74, 75, 76], variational Monte Carlo [77] techniques, and quantum computing variational approaches [78, 79, 80]. Moreover, orbital optimizations are common practice on SCI-based approaches as well [81]. More generally, orbital optimizations make variational approaches invariant under orbital rotations. For notation clarity we replace in this section the notation of the two-body integral $(pr|qs)$ by the abbreviation $h_{pqrs}$.
The implementation of the orbital optimizations in conjunction with the HPC quantum estimator (with or without configuration recovery) follows the procedure presented in Ref. [77]. Orbital rotations consist of the application of the similarity transformation:
$\hat{\tilde{H}}=\hat{U}^{\dagger}(\kappa)\hat{H}\hat{U}(\kappa),$ (S32)
where
In this work, we choose the matrix that parametrizes the rotation to be real: $\kappa\in\mathbb{R}^{N_{\mathrm{MO}}\times N_{\mathrm{MO}}}$. To enforce unitarity in $\hat{U}(\kappa)$ we require that $\kappa_{pq}=-\kappa_{qp}$. According to Thouless’s Theorem [82], the action of the similarity transformation defined by $\hat{U}(\kappa)$ transforms the creation operators according to:
using Einstein’s summation convention.
A variational procedure is used to search for the single-particle basis that yields the optimal description of the ground state, given the collection of all possible $K_{\textrm{max}}$ bare variational trial states $\left|\psi_{\Theta}^{(k)}\right\rangle$. The variational procedure is similar to the method presented in Sec. II.1.3. In this setting, $\left|\psi_{\Theta}^{(k)}\right\rangle$ is the SCI ansatz given as the linear combination of the $d$ determinants in $\mathcal{S}^{(k)}$ (see Eq. (S10)). The set of variational parameters $\Theta$ is composed of the circuit parameters $\theta$ defining $\left|\Psi_{\theta}\right\rangle$ together with the wavefunction amplitudes $c_{\bf{x}}^{(k)}$ in the subspace defined by the SCI configurations. Each variational state is dressed by the same single-particle orbital rotation
Gradient descent and its variants can be used to minimize both $\kappa$ and $\theta$ in $\mathcal{E}(\kappa,\Theta)$, while the optimization of $c_{{\bf{x}}}^{(k)}$ is carried out by the diagonalization procedure (see Sec. II.1.1). Note that the sum in the average contains double-combinatorially many terms. Consequently, a Monte-Carlo based estimator is used to obtained an unbiased estimate (see Sec. II.1.3). Gradients with respect to $\theta$ have the same expression as in Eq. (S29) replacing $\hat{H}$ by $\hat{\tilde{H}}$. Gradients with respect to the orbital rotations can be computed by the contraction of the bare one- and two-body reduced density matrices (1- and 2-RDMs) with the gradients of the one- and -two body integrals with respect to $\kappa_{pq}$ in Eq. (S38):
and the bare 1-RDMs and 2-RDMs are defined as
The gradients of the integrals with respect to the rotation parameters are computed using the automatic differentiation (AD) tools of the software package Jax [83]. Since the evaluation of the gradients with respect to $\kappa$ requires the evaluation of an intractable sum, we use an unbiased Monte-Carlo estimator for its evaluation:
See Sec. V.3 to see the effect of orbital optimizations in the accuracy of the HPC quantum estimator (with configuration recovery). The section shows results run on measurement outcomes obtained from the Heron device, to study the dissociation of N2 at cc-pVDZ level of theory.
II.2 Wavefunction concentration and accuracy
II.2.1 Sample complexity
@@
II.2.2 Assessing wavefunction concentration
@@
II.2.3 A numerical study of the accuracy: Hubbard model
To illustrate the dependence of the accuracy of the SCI-based eigensolver as a function of the ground-state wavefunction concentration we consider the $N_{\mathrm{MO}}$-site Hubbard model in a fully-connected lattice with random hopping amplitudes:
@@ II.2.3 A numerical study of the accuracy: Hubbard model
The hopping amplitudes $t_{pq}=t_{qp}$ are independent random variables drawn from the Gaussian distribution with mean 00 and standard deviation $1$, i.e. $\overline{t_{pq}}=0$ and $\overline{t^{2}_{pq}}=t^{2}=1$. The results presented in this section are obtained from the average of results obtained from twenty disorder realizations. In the infinite-volume, this model is known to possess self-averaging properties [84]. However, on finite-size systems, there are sample-to-sample fluctuations. The numerical study in this section does not consider the effect of noise, so we run the estimator without configuration recovery. For the discussion that follows, we span the wavefunction amplitudes in terms of electronic configurations in the localized basis, i.e. the basis used to define the Hamiltonian in Eq. (S56), contrary to the molecular systems considered where the reference basis is that of Hartree-Fock orbitals.
We choose the Hubbard model for this analysis because the ground-state properties, and the wavefunction concentration, depend on a single parameter: $U$. Furthermore, the study of the Hubbard model with various flavors of Configuration Interaction approaches has been considered by previous works [85, 86, 87, 88]. The onsite Hubbard repulsion $U$ controls the nature of the ground state of the system. In the $N_{\mathrm{MO}}\rightarrow\infty$ limit, and for $0\leq U<U_{c}$, the ground state is a Fermi liquid metal. For $U>U_{c}$, the ground state is a Mott insulator. The metal to insulator transition in this case is of second order [89, 90]. Dynamical Mean Field Theory studies report a transition point of $U_{c}=5.82...$ [91]. The transition on finite-size systems is harder to identify [92]. The change in the nature of the ground state is reflected in the concentration of the wavefunction amplitudes. In the Mott phase, electrons tend to localize and there is a preference for configurations with low double occupancy [92]. For $U=0$ and small $U$, and given the disordered nature of the model, it is expected that all electronic configurations will be equivalent, especially after considering a number of disorder realizations. It is therefore expected that the eigenstate solver will show better accuracy for larger values of $U$.
Since the goal of this numerical experiment is to isolate the performance of the SCI eigensolver as a function of the wavefunction concentration, the electronic configurations used to run the SCI eigenstate solver are the $d=10^{3}$ configurations of highest amplitude in the exact ground state. This choice removes any effects in the accuracy coming from the choice of the quantum circuit generating the configurations . This procedure is equivalent to the QSCI proposal [30], where the configurations are sampled from the exact ground-state wavefunction. As a measure of concentration, we consider the entropy of the distribution defined by the wavefunction amplitudes:
@@
To measure the accuracy in the ground state we consider the relative error in the ground-state energy, a common metric in variational many-body calculations. The relative error is refined as: $\left|E_{\textrm{SCI}}-E_{\textrm{exact}}\right|/E_{\textrm{exact}}$. For the numerical experiments we consider $N_{\mathrm{MO}}=10$, and 31 equally-spaced values of $U$, between $U=1$ and $U=16$, both included.
@@
Fig. S3 (a) shows the relative error in the ground-state energy obtained by SCI as a function of the entropy of the distribution of the wavefunction amplitudes. The more concentrated the ground-state wavefunction (lower entropy), the lower the relative error becomes. As expected, the SCI-based eigensolver performs better for large $U$ deep in the Mott phase. For small $U$ in the metallic phase, the SCI trial wavefunction with $d=10^{3}$ does not have enough expressive power to accurately represent the ground state of the system, whose amplitudes anti-concentrate.

### v3

@@ Results
  3. 3.
    We extend the unitaries to the following matrices,
    $\tilde{U}_{y,pr}=\delta_{pa}\delta_{ri}U_{ai,y}\;,$ (23)
    i.e. matrices where only the occupied/unoccupied block is non-zero.
  4. 4.
    We define the Hermitian operators
    $X_{\pm,y}=\frac{1\mp i}{2}\left(\tilde{U}_{y}\pm i\tilde{U}_{y}^{T}\right)$ (24)
    and their eigenpairs
    $X_{\pm,y}V_{\pm,y}=g_{\pm,y}V_{\pm,y}~{}.$ (25)
  5. 5.
    We define the operators
    $\begin{split}\left(J_{2y}\right)^{\sigma\tau}_{pr}&=\tau_{y}\left(g_{+,y}\right)_{p}\left(g_{+,y}\right)_{r}\\ \left(K_{2y}\right)^{\sigma}_{pr}&=\log(V_{+,y})_{pr}\\ \left(J_{2y+1}\right)^{\sigma\tau}_{pr}&=-\tau_{y}\left(g_{-,y}\right)_{p}\left(g_{-,y}\right)_{r}\\ \left(K_{2y+1}\right)^{\sigma}_{pr}&=\log(V_{-,y})_{pr}~{}.\\ \end{split}$ (26)
  6. 6.
    We retain the first $L$ matrices $J$, $K$. This allows us to refine a non-local UCJ wavefunction  (?, ?).
  7. 7.
    $\begin{split}&\left(J_{\mu}\right)^{\alpha\alpha}_{p,p+1}\;,\;p=0\dots N_{\mathrm{MO}}-2\\ &\left(J_{\mu}\right)^{\beta\beta}_{p,p+1}\;,\;p=0\dots N_{\mathrm{MO}}-2\\ &\left(J_{\mu}\right)^{\alpha\beta}_{p,p}\;,\;p=0\dots N_{\mathrm{MO}}-1\;,\;p\%4=0~{}.\\ \end{split}$ (27)
Mapping to heavy-hex processors
The choice of retaining the elements $(J_{\mu})^{\alpha\beta}_{p,p}$ with $p=0\dots N_{\mathrm{MO}}-1\;,\;p\%4=0$ when implementing LUCJ on a heavy-hex processor has an important technical motivation: on such devices, assuming a number of qubits greatly exceeding $N_{\mathrm{MO}}$, spin-up and spin-down orbitals can be mapped on two segments of adjacent qubits forming a “zig-zag” pattern and connected through an auxiliary qubit for $p=0,4,8,\dots$ as shown in the three rightmost panels of Fig. 4. Such a qubit layout allows implementing LUCJ with a minimal overhead of $\mathsf{SWAP}$ gates (two per auxiliary qubit and layer of LUCJ). However, on current processors with up to 133 qubits, for $N_{\mathrm{MO}}>21$ one cannot couple $N_{\mathrm{MO}}/4$ spin-orbitals with opposite spins through auxiliary qubits without incurring a significant overhead of $\mathsf{SWAP}$ gates, for the simple reason that a chain of 22 or more qubits is longer than the “diagonal” of the processor. In such a situation, as shown in the rightmost panel of Fig. 4, the segments on which spin-up and spin-down qubits are mapped form two “tails” that are not connected by auxiliary qubits. This fact has two implications: (i) no more than 6 spin-orbitals with opposite spins can be coupled through auxiliary qubits, and (ii) if one retains the elements $(J_{\mu})^{\alpha\beta}_{p,p}$ with $p=0\dots N_{\mathrm{MO}}-1\;,\;p\%4=0$ and $p\leq 16$, the largest elements of $J_{\mu}$ may be discarded, yielding a lower-accuracy wavefunction. The first problem is a fundamental one, that can only be resolved with a substantially different mapping of fermionic degrees of freedom onto qubits and/or through the availability of larger processors. The second problem, on the other hand, has a simple solution, that we now describe. First, for any pair of spatial orbitals $p,r=0\dots N_{\mathrm{MO}}-1$, the orbital rotation
$\hat{S}_{pr}=e^{-i\frac{\pi}{2}\sum_{\sigma}\left(\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{r\sigma}+\hat{a}^{\dagger}_{r\sigma}\hat{a}^{\phantom{\dagger}}_{p\sigma}\right)}$ (28)
$\hat{S}^{\dagger}_{pr}\left(\sum_{qs,\tau}M_{qs}\hat{a}^{\dagger}_{q\tau}\hat{a}^{\phantom{\dagger}}_{s\tau}\right)\hat{S}_{pr}=\sum_{qs,\tau}M^{\prime}_{qs}\hat{a}^{\dagger}_{q\tau}\hat{a}^{\phantom{\dagger}}_{s\tau}\;,\;M^{\prime}_{qs}=M_{S_{pr}(q),S_{pr}(s)}\;.$ (29)
Three immediate implications of this fact are:
  • $\hat{S}^{\dagger}_{pr}\hat{n}_{q\tau}\hat{S}_{pr}=\hat{n}_{S_{pr}(q)\tau}$
    $\hat{S}^{\dagger}_{pr}\hat{J}_{\mu}\hat{S}_{pr}=\sum_{qs,\tau}(J^{\prime}_{\mu})^{\sigma\tau}_{q,s}\hat{n}_{q\sigma}\hat{n}_{s\tau}=\hat{J}^{\prime}_{\mu}\;,\;(J^{\prime}_{\mu})^{\sigma\tau}_{q,s}=({J}_{\mu})^{\sigma\tau}_{S_{pr}(q),S_{pr}(s)}\;.$ (30)
$e^{-\hat{K}_{S}}\hat{J}_{\mu}e^{\hat{K}_{S}}=\hat{J}^{\prime}_{\mu}$ (31)
$e^{\hat{K}_{\mu}}e^{i\hat{J}_{\mu}}e^{-\hat{K}_{\mu}}=e^{\hat{K}_{\mu}}e^{\hat{K}_{S}}e^{i\hat{J}^{\prime}_{\mu}}e^{-\hat{K}_{S}}e^{-\hat{K}_{\mu}}=e^{\hat{K}^{\prime}_{\mu}}e^{i\hat{J}^{\prime}_{\mu}}e^{-\hat{K}^{\prime}_{\mu}}$ (32)
to obtain a UCJ operator with transformed orbital rotations $e^{-\hat{K}^{\prime}_{\mu}}$ and an opposite-spin density-density interaction whose largest elements in absolute value act on spatial orbitals $p=0,4,8,\dots,4\ell$. Sparsification of $(J^{\prime}_{\mu})^{\alpha\beta}_{q,s}$ then leads to retaining the $l$ dominant opposite-spin density-density interaction terms (as many as allowed by the size and topology of the available heavy-hex processor) with a minimal overhead of $\mathrm{SWAP}$ gates.
Acknowledgments
We thank the IBM Quantum Service and Data team for help with the workflow execution. The authors acknowledge feedback and insightful conversations with Alberto Baiardi, Sergey Bravyi, Giuseppe Carleo, Garnet Kin-Lic Chan, Antonio Córcoles, Oliver Dial, Jay Gambetta, Artur Izmaylov, Caleb Johnson, Yukio Kawashima, David Kremer, Isaac Lauer, Peter Love, Guglielmo Mazzola, Takahito Nakajima, Paul Nation, Hanhee Paik, Emily Pritchett, Max Rossmannek, Paul Schweigert, James E.T. Smith, Miles Stoudenmire, Nobuyuki Yoshioka, Andrew Wack, and Christa Zoufal. A part of this work is supported by project JPNP20017, funded by the New Energy and Industrial Technology Development Organization (NEDO), Japan, and by the RIKEN TRIP initiative (RIKEN Quantum).
Author contributions:
Design of the workflow and experiments: J. R.-M., M. M., W. K., K. S., M. C. T., A. M. Implementation of the workflow: J. R.-M., M. M., A. J.-A., S. M., S. S., T. S., I. S., R.-Y. S., K. J. S. Tuning and calibration of the Heron processor: H.H. Execution of the quantum part of the workflow on Heron: M. M., P. J., M. T. Execution of the classical part of the workflow on Fugaku: T. S., R.-Y. S., S. Y. Numerical benchmarks: J. R.-M., M.M., T. S., K. J. S. Analytical derivations: J. R.-M., M.M., W. K., K. S., M. C. T., A. M. All authors contributed to the manuscript writing and data analysis.
Data and materials availability:
All data supporting the findings is available in the main text and Supplementary materials. The code to reproduce all results reported is available as the open-source package qiskit-addon-sqd  (?).
Supplementary materials
Supplementary Text
Figs. S1 to S22
Tables S1 to S2
References (7-0)
Supplementary Materials for
Chemistry Beyond the Scale of Exact Diagonalization on a Quantum-Centric Supercomputer
Javier Robledo-Moreno1†, Mario Motta1∗, Holger Haas1, Ali Javadi-Abhari1,Petar Jurcevic1,William Kirby2,Simon Martiel3,Kunal Sharma1,Sandeep Sharma4,Tomonori Shirakawa5,6,7,Iskandar Sitdikov1,Rong-Yang Sun5,6,7,Kevin J. Sung1,Maika Takita1,Minh C. Tran2,Seiji Yunoki5,6,7,8,Antonio Mezzacapo1⟂.1IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598, USA.2IBM Quantum, IBM Research Cambridge, Cambridge, MA 02142, USA.3IBM Quantum, IBM France Lab, Orsay, France.4Department of Chemistry, University of Colorado, Boulder, CO 80302, USA.5Computational Materials Science Research Team, RIKEN Center for Computational Science (R-CCS), Kobe, Hyogo, 650-0047, Japan.6Quantum Computational Science Research Team, RIKEN Center for Quantum Computing (RQC),Wako, Saitama, 351-0198, Japan.7RIKEN Interdisciplinary Theoretical and Mathematical Sciences Program (iTHEMS), Wako, Saitama 351-0198, Japan. 8RIKEN Center for Emergent Matter Science (CEMS), Wako, Saitama 351-0198, Japan.Corresponding author. Email: j.robledomoreno@ibm.comCorresponding author. Email: mario.motta@ibm.comCorresponding author. Email: mezzacapo@ibm.com
This PDF file includes:
Supplementary Text
Figures S1 to S22
Tables S1 to S2
1 Convergence properties of the components of sample-based quantum diagonalization
1.1 Wavefunction concentration and accuracy
1.1.1 Sample complexity
@@
1.1.2 Assessing wavefunction concentration
@@
1.1.3 A numerical study of the accuracy: Hubbard model
To illustrate the dependence of the accuracy of the Selected Configuration Interaction (SCI)-based eigensolver as a function of the ground-state wavefunction concentration we consider the $N_{\mathrm{MO}}$-site Hubbard model in a fully-connected lattice with random hopping amplitudes:
@@ 1.1.3 A numerical study of the accuracy: Hubbard model
The hopping amplitudes $t_{pq}=t_{qp}$ are independent random variables drawn from the Gaussian distribution with mean 00 and standard deviation $1$, i.e. $\overline{t_{pq}}=0$ and $\overline{t^{2}_{pq}}=t^{2}=1$. The results presented in this section are obtained from the average of results obtained from twenty disorder realizations. In the infinite-volume, this model is known to possess self-averaging properties  (?). However, on finite-size systems, there are sample-to-sample fluctuations. The numerical study in this section does not consider the effect of noise, so we run the estimator without configuration recovery. For the discussion that follows, we span the wavefunction amplitudes in terms of electronic configurations in the localized basis, i.e. the basis used to define the Hamiltonian in Eq. (S14), contrary to the molecular systems considered where the reference basis is that of Hartree-Fock orbitals.
We choose the Hubbard model for this analysis because the ground-state properties, and the wavefunction concentration, depend on a single parameter: $U$. Furthermore, the study of the Hubbard model with various flavors of Configuration Interaction approaches has been considered by previous works  (?, ?, ?, ?). The onsite Hubbard repulsion $U$ controls the nature of the ground state of the system. In the $N_{\mathrm{MO}}\rightarrow\infty$ limit, and for $0\leq U<U_{c}$, the ground state is a Fermi liquid metal. For $U>U_{c}$, the ground state is a Mott insulator. The metal to insulator transition in this case is of second order  (?, ?). Dynamical Mean Field Theory studies report a transition point of $U_{c}=5.82...$  (?). The transition on finite-size systems is harder to identify  (?). The change in the nature of the ground state is reflected in the concentration of the wavefunction amplitudes. In the Mott phase, electrons tend to localize and there is a preference for configurations with low double occupancy  (?). For $U=0$ and small $U$, and given the disordered nature of the model, it is expected that all electronic configurations will be equivalent, especially after considering a number of disorder realizations. It is therefore expected that the eigenstate solver will show better accuracy for larger values of $U$.
Since the goal of this numerical experiment is to isolate the performance of SQD as a function of the wavefunction concentration, the electronic configurations used to run the SCI eigenstate solver are the $d=10^{3}$ configurations of highest amplitude in the exact ground state. This choice removes any effects in the accuracy coming from the choice of the quantum circuit generating the configurations. This procedure is equivalent to the QSCI proposal  (?), where the configurations are sampled from the exact ground-state wavefunction. As a measure of concentration, we consider the entropy of the distribution defined by the wavefunction amplitudes:
@@
To measure the accuracy in the ground state we consider the relative error in the ground-state energy, a common metric in variational many-body calculations. The relative error is refined as: $\left|E_{\textrm{SQD}}-E_{\textrm{exact}}\right|/E_{\textrm{exact}}$. For the numerical experiments we consider $N_{\mathrm{MO}}=10$, and 31 equally-spaced values of $U$, between $U=1$ and $U=16$, both included.
@@
Fig. S1 (a) shows the relative error in the ground-state energy obtained by SQD as a function of the entropy of the distribution of the wavefunction amplitudes. The more concentrated the ground-state wavefunction (lower entropy), the lower the relative error becomes. As expected, SQD performs better for large $U$ deep in the Mott phase. For small $U$ in the metallic phase, the SCI trial wavefunction with $d=10^{3}$ does not have enough expressive power to accurately represent the ground state of the system, whose amplitudes anti-concentrate.